Value of (a – c) [(a – b)2 + (b – c)2 – (a – b) (b – c)] + (c – a)3 is equal to

Grade:10

1 Answers

Shailendra Kumar Sharma
6 years ago
(a – c) [(a – b)2 + (b – c)2 – (a – b) (b – c)] + (c – a)3
Take (a – c) [(a – b)2 + (b – c)2 – (a – b) (b – c)] 
Take a-b =p and b-c=q then
(p+q)(p2-pq+q2) which is equal to (p+q)3
So 3 =(a-c)3 ={(a-c)+(b-c)}(a – c) [(a – b)2 + (b – c)2 – (a – b) (b – c)] 
So (a – c) [(a – b)2 + (b – c)2 – (a – b) (b – c)] + (c – a)3 =  =03+(c-a)(a-c)